# Grover Speedup from Many Forms of the Zeno Effect

### v3

@@ C.1 Purity arguments for measurements and decoherence
From these equations it follows that,
@@
In the pure case in which $\mathrm{Tr}\left[\rho^{2}\right]=1$, Eqn (61) becomes $p_{m}=(1\pm 1)/2$, so $p_{m}$ is just 00 or $1$. This makes intuitive sense: a pure state could be the state $|m\rangle$ (corresponding to the value of $1$), or any other state (corresponding to the value of 00), but any non-trivial sum of the two will no longer be pure [^4] Recall that the probability to be measured in $|m\rangle$ is not simply $p_{m}$, but $p_{m}+p_{\psi}\left|\left\langle m\middle|\psi\right\rangle\right|^{2}$ so $p_{m}$ is necessarily zero for pure states with non-unity probability to be found in the state $|m\rangle$. On the other extreme, a maximally mixed state, $\mathrm{Tr}\left[\rho^{2}\right]=1/2$, $p_{m}=1/2$. For intermediate values, this formula gives a non-zero lower bound on the value of $p_{m}$, since the negative branch will yield a smaller $p_{m}$ value for $11/2\leq\mathrm{Tr}\left[\rho^{2}\right]<$ 1, it follows that $p_{m}\geq\frac{1}{2}\left(1-\sqrt{1-2\left(1-\mathrm{Tr}\left[\rho^{2}\right]\right)}\right)$. Finally this can be used to bound the total probability of measuring the marked state,
@@
C.2 Ground states preserved by Zeno effects
As discussed in the main text, the previous argument does not strictly apply to the setting of a Zeno effect since the Zeno effect describes repeated measurements which maintain a pure state. Since rotation of the underlying state space in an avoided crossing is different from the traditional setting of “freezing” time evolution, it is worth reviewing how to derive the fact that rapid removal of off-diagonal density matrix elements in the instantaneous energy eigenbasis can maintain the ground state.
@@ C.2 Ground states preserved by Zeno effects
From this formula, we can immediately see that if we start in the ground state $\rho_{gg}=1$ and $\rho_{ee}=\rho_{ge}=0$ and a phase is introduced to $\rho_{ge}$ at each step, than the sign of $\mathrm{Re}[\rho_{ge}]$ will eventually flip causing, the flow of amplitude to reverse, flowing instead from the excited state to the ground state, this is the underlying mechanism behind both adiabatic operation and the multi-stage quantum walk manifestation of the Zeno effect. An extreme manifestation occurs when phase flipping occurs, in which case the flow of probability reverses direction in a single step, returning the lost probability to the ground state.
@@
if we now start in the ground state $\rho_{gg}=1$ and $\rho_{ee}=\rho_{ge}=0$ and perform successive operations where $\rho_{ee}\rightarrow\rho_{ee}+\delta^{2}_{j}\rho_{gg}$, we find that $\rho_{ee}=\sum_{j}\delta^{2}_{j}\approx 0$. This illustrates the presence of a Zeno effect which maintains the instantaneous ground state since $\rho_{ee}+\rho_{gg}=1$ by definition. We further note that removal of the excited state at each step does not fundamentally change this picture, it just adds an additional step where $|e\rangle\rightarrow|d\rangle$, and therefore similarly $\rho_{dd}=\sum_{j}\delta^{2}_{j}\approx 0$. These arguments apply equally to rapid continuous dissipation or decoherence.
Appendix D Derivation of master equations
D.1 Derivation of equation 28
The first stage of the derivation is to consider the average case of an instance of equation 22 between two $\tilde{X}$ operations. We will assume that this is operation is occurring after many flips have occurred so there is a 50% chance that an odd number of $\tilde{X}$ operations have been applied and a 50% chance an even number have been applied. During this time period the evolution will be described as
$U_{\mp}(q,j_{0})=\mathcal{T}\prod^{j_{0}+m/q}_{j=j_{0}}\exp\left(\frac{\mp it_{\mathrm{scale}}\sqrt{\gamma_{j}+1}}{m}\tilde{Z}\right)=$
$\mathcal{T}\prod^{j_{0}+m/q}_{j=j_{0}}\exp\left(\frac{\mp it_{\mathrm{scale}}\sqrt{f(\tau_{j})+1}}{m}\tilde{Z}\right)$ (72)
@@ D.1 Derivation of equation 28
$U_{\mp}(q,\tau)\approx\mathcal{T}\prod^{j_{0}+m/q}_{j=j_{0}}\exp\left(\frac{\mp it_{\mathrm{scale}}\sqrt{f(\tau)+1}}{m}\tilde{Z}\right)=$
$\left[\exp\left(\frac{\mp it_{\mathrm{scale}}\sqrt{f(\tau)+1}}{m}\tilde{Z}\right)\right]^{m/q}=\exp\left(\frac{\mp it_{\mathrm{scale}}\sqrt{f(\tau)+1}}{q}\tilde{Z}\right).$ (73)
@@
$\Phi=-\frac{t_{\mathrm{scale}}\sqrt{f(\tau)+1}}{q}.$ (75)
@@
$\frac{\Delta\rho}{\Delta t}\approx\frac{\Phi^{2}q}{2t_{\mathrm{scale}}}\left(\tilde{Z}\rho\tilde{Z}-\rho\right)=t_{\mathrm{scale}}\frac{f(\tau)+1}{2q}\left(\tilde{Z}\rho\tilde{Z}-\rho\right)=\kappa\left(\tilde{Z}\rho\tilde{Z}-\rho\right),$ (79)
@@
which is equation 28. An astute reader may be concerned that this derivation assumes that a large number of $\tilde{X}$ operations have already occurred. Recall however that at the initial points of the evolution the ground and excited states will be unchanged and to a very good approximation $|\tilde{\omega}\rangle$ and $|m\rangle$, with the system starting in $|\tilde{\omega}\rangle$. In this case the relative phase between the ground and first excited state is fairly irrelevant, allowing time for randomisation in the phase directions to build up before the approach to the crossing.
D.2 Derivation of equation 40
@@
for notational convenience. We now consider the effect of taking a projective measurement to determine if the auxilliary system is in the $|d\rangle$ state and abort the algorithm if it is. For notational simplicity, we consider an aborted algorithm as placing the entire system (as opposed to just the auxilliary) into a computationally useless $|d\rangle$ state. By the Born measurement rule, the probability of aborting the algorithm is $\sin^{2}\left(\Phi\right)\approx\Phi^{2}$. Since the system can only be put into the $|d\rangle$ state if it started in the $|e\rangle$ state, the overall effect on the density matrix is,
@@
Appendix E Equation 36 in the density matrix formalism and an alternative model for decoherence
@@
is the mathematical operation which sends the system to the $|d\rangle$ state if the additional qubit is in the $|1\rangle$ state and leaves it unchanged otherwise. The remaining degrees of freedom are left for mathematical convenience, for our purposes any system where the last degree of freedom is in the $|d\rangle$ state is considered computationally useless [^5] To avoid this construction one can instead construct a superoperator which acts on $\rho$ and separately projects each possible state of the system into a global $|d\rangle\!\langle d|$ state, but we have elected to maintain these extra degrees of freedom to simplify the expression.. Any operation is considered to act as the identity on the $|d\rangle$ state in other words $U\rightarrow U\oplus\mathbb{1}_{1}$ for all of the unitary operations within equation 89 where the one-dimensional identity acts on the destroyed state which has been added to the Hilbert space.

### v4

@@ D.1 Purity arguments for measurements and decoherence
From these equations, it follows that,
@@
In the pure case in which $\mathrm{Tr}\left[\rho^{2}\right]=1$, Eqn (70) becomes $p_{m}=(1\pm 1)/2$, so $p_{m}$ is just 00 or $1$. This makes intuitive sense: a pure state could be the state $|m\rangle$ (corresponding to the value of $1$), or any other state (corresponding to the value of 00), but any non-trivial sum of the two will no longer be pure [^6] Recall that the probability to be measured in $|m\rangle$ is not simply $p_{m}$, but $p_{m}+p_{\psi}\left|\left\langle m\middle|\psi\right\rangle\right|^{2}$ so $p_{m}$ is necessarily zero for pure states with non-unity probability to be found in the state $|m\rangle$. On the other extreme, a maximally mixed state, $\mathrm{Tr}\left[\rho^{2}\right]=1/2$, $p_{m}=1/2$. For intermediate values, this formula gives a non-zero lower bound on the value of $p_{m}$, since the negative branch will yield a smaller $p_{m}$ value for $11/2\leq\mathrm{Tr}\left[\rho^{2}\right]<$ 1, it follows that $p_{m}\geq\frac{1}{2}\left(1-\sqrt{1-2\left(1-\mathrm{Tr}\left[\rho^{2}\right]\right)}\right)$. Finally, this can be used to bound the total probability of measuring the marked state,
@@
D.2 Ground states preserved by Zeno effects
As discussed in the main text, the previous argument does not strictly apply to the setting of a Zeno effect, since the Zeno effect describes repeated measurements which maintain a pure state. Since rotation of the underlying state space in an avoided crossing is different from the traditional setting of “freezing” time evolution, it is worth reviewing how to derive the fact that rapid removal of off-diagonal density matrix elements in the instantaneous energy eigenbasis can maintain the ground state.
@@ D.2 Ground states preserved by Zeno effects
From this formula, we can immediately see that if we start in the ground state $\rho_{gg}=1$ and $\rho_{ee}=\rho_{ge}=0$ and a phase is introduced to $\rho_{ge}$ at each step, than the sign of $\mathrm{Re}[\rho_{ge}]$ will eventually flip causing, the flow of amplitude to reverse, flowing instead from the excited state to the ground state, this is the underlying mechanism behind both adiabatic operation and the multi-stage quantum walk manifestation of the Zeno effect. An extreme manifestation occurs when phase flipping is studied, in which case the flow of probability reverses direction in a single step, returning the lost probability to the ground state.
@@
if we now start in the ground state $\rho_{gg}=1$ and $\rho_{ee}=\rho_{ge}=0$ and perform successive operations where $\rho_{ee}\rightarrow\rho_{ee}+\delta^{2}_{j}\rho_{gg}$, we find that $\rho_{ee}=\sum_{j}\delta^{2}_{j}\approx 0$. This illustrates the presence of a Zeno effect, which maintains the instantaneous ground state since $\rho_{ee}+\rho_{gg}=1$ by definition. We further note that removal of the excited state at each step does not fundamentally change this picture, it just adds an additional step where $|e\rangle\rightarrow|d\rangle$, and therefore similarly $\rho_{dd}=\sum_{j}\delta^{2}_{j}\approx 0$. These arguments apply equally to rapid continuous dissipation or decoherence.
Appendix E Derivation of master equations
E.1 Derivation of equation 28
The first stage of the derivation is to consider the average case of an instance of equation 22 between two $\tilde{X}$ operations. We will assume that this is operation is occurring after many flips have occurred, so there is a 50% chance that an odd number of $\tilde{X}$ operations have been applied and a 50% chance an even number have been applied. During this time period the evolution will be described as
$U_{\mp}(q,j_{0})=\mathcal{T}\prod^{j_{0}+m/q}_{j=j_{0}}\exp\left(\frac{\mp it_{\mathrm{scale}}\sqrt{\gamma^{2}_{j}+1}}{m}\tilde{Z}\right)=$
$\mathcal{T}\prod^{j_{0}+m/q}_{j=j_{0}}\exp\left(\frac{\mp it_{\mathrm{scale}}\sqrt{f^{2}(\tau_{j})+1}}{m}\tilde{Z}\right)$ (81)
@@ E.1 Derivation of equation 28
$U_{\mp}(q,\tau)\approx\mathcal{T}\prod^{j_{0}+m/q}_{j=j_{0}}\exp\left(\frac{\mp it_{\mathrm{scale}}\sqrt{f^{2}(\tau)+1}}{m}\tilde{Z}\right)=$
$\left[\exp\left(\frac{\mp it_{\mathrm{scale}}\sqrt{f^{2}(\tau)+1}}{m}\tilde{Z}\right)\right]^{m/q}=\exp\left(\frac{\mp it_{\mathrm{scale}}\sqrt{f(^{2}\tau)+1}}{q}\tilde{Z}\right).$ (82)
@@
$\Phi=-\frac{t_{\mathrm{scale}}\sqrt{f^{2}(\tau)+1}}{q}.$ (84)
@@
$\frac{\Delta\rho}{\Delta t}\approx\frac{\Phi^{2}q}{2t_{\mathrm{scale}}}\left(\tilde{Z}\rho\tilde{Z}-\rho\right)=t_{\mathrm{scale}}\frac{f^{2}(\tau)+1}{2q}\left(\tilde{Z}\rho\tilde{Z}-\rho\right)=\kappa\left(\tilde{Z}\rho\tilde{Z}-\rho\right),$ (88)
@@
which is equation 28. An astute reader may be concerned that this derivation assumes that many $\tilde{X}$ operations have already occurred. Recall however that at the initial points of the evolution the ground and excited states will be unchanged and to a very good approximation $|\tilde{\omega}\rangle$ and $|m\rangle$, with the system starting in $|\tilde{\omega}\rangle$. In this case, the relative phase between the ground and first excited state is fairly irrelevant, allowing time for randomisation in the phase directions to build up before the approach to the crossing.
E.2 Derivation of equation 40
@@
for notational convenience. We now consider the effect of taking a projective measurement to determine if the auxilliary system is in the $|d\rangle$ state, and abort the algorithm if it is. For notational simplicity, we consider an aborted algorithm as placing the entire system (as opposed to just the auxilliary) into a computationally useless $|d\rangle$ state. By the Born measurement rule, the probability of aborting the algorithm is $\sin^{2}\left(\Phi\right)\approx\Phi^{2}$. Since the system can only be put into the $|d\rangle$ state if it started in the $|e\rangle$ state, the overall effect on the density matrix is,
@@
Appendix F Equation 36 in the density matrix formalism and an alternative model for decoherence
@@
is the mathematical operation which sends the system to the $|d\rangle$ state if the additional qubit is in the $|1\rangle$ state and leaves it unchanged otherwise. The remaining degrees of freedom are left for mathematical convenience, for our purposes any system where the last degree of freedom is in the $|d\rangle$ state is considered computationally useless [^7] To avoid this construction one can instead construct a superoperator which acts on $\rho$ and separately projects each possible state of the system into a global $|d\rangle\!\langle d|$ state, but we have elected to maintain these extra degrees of freedom to simplify the expression.. Any operation is considered to act as the identity on the $|d\rangle$ state, in other words $U\rightarrow U\oplus\mathbb{1}_{1}$ for all of the unitary operations within equation 98 where the one-dimensional identity acts on the destroyed state which has been added to the Hilbert space.